OpenStudy (anonymous):

Mrs. Stephens has a bag of candy. The ratio of peppermints to chocolates is 5:3, and the ratio of peppermints to gummies is 3:4. What is the ratio of chocolates to gummies? Express your answer as a common fraction.

OpenStudy (anonymous):

I'm pretty sure 9/20, not 100% though, a little sleepy.

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

@taluvsdawgs Let the number of peppermints be represented by p, the number of chocolates c, gummies g. Given the ratios in the question, we know that: p/c = 5/3 and p/g = 3/4 We want to know what (c/g) is. Proceed using simple algebra: p = (5/3) * c p = (3/4) * g Using a systems of equations "equal values" style approach, we get: (5/3) * c = (3/4) * g Then: (5/3) * (c/g) = 3/4 Multiply by the reciprocal of 5/3 to isolate c/g: c/g = (3/4) * (3/5) c/g = 9/20 The ratio is 9:20, already in lowest terms. Let me know if any of this is unclear.

OpenStudy (anonymous):

it is clear thanks a lot